OpenStudy (anonymous):

Write 32x^5y^5 with only one exponent. Use parentheses?

OpenStudy (callisto):

2^5 =32 therefore 32x^5y^5 = 2^5 x^5 y^5 Can you do it from here?

OpenStudy (anonymous):

2xy^15

OpenStudy (callisto):

nope... \[a^x b^x = (ab)^x\] From this, can you try again?

OpenStudy (anonymous):

32x^25y

OpenStudy (callisto):

Hmm... \[a^5b^5c^5 = (abc)^5\] Can you try again?

OpenStudy (callisto):

Hint: a =2, b=x, c=y ...

OpenStudy (anonymous):

32xy

OpenStudy (callisto):

nope....

OpenStudy (callisto):

32x^5 y^5 = 2^5 x^5 y^5

OpenStudy (anonymous):

ok 2^5 is 32 right

OpenStudy (callisto):

yes

OpenStudy (anonymous):

32xy^10

OpenStudy (callisto):

no you cannot do it like that \[x^5y^5 \neq xy^{10}\]

OpenStudy (callisto):

\[x^5y^5 = (xy)^5\]

OpenStudy (anonymous):

32(xy)^5

OpenStudy (callisto):

Perhaps, but personally i'd convert the 32 too... Since 32 =2^5 \[2^5 x^5 y^5 = (2xy)^5\]

OpenStudy (anonymous):

see i new that but when i put it like that on my test it was wrong but i'll put it again and have my teacher chek it

OpenStudy (callisto):

Okay, post the result once you've got it, I also want to check it too!